Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-x-3y=.20
5x+3y=40

Answer by lawdej(58)

here is the way
-x-3y=.20
5x+3y=40
take one of the equations as you like
-x-3y=.20
x=-3y-.20 substitute this into 5x+3y=40
5(-3y-.20)+3y=40
-15y-1+3y=40
-12y=41
y=-41/12
to find value of x substitute y=-41/12 in either equations as you like
x=-3y-.20=-3(-41/12)-.20
x=(123-2.4)/14=120.6/12
x=120.6/14